3 ms·
I’ve seen LLM catch and fix their own mistakes and literally tell me they were wrong and that they are fixing their self made wrong mistake. This analogy is the
by ninetyninenine 1y ago
I’ve seen LLM catch and fix their own mistakes and literally tell me they were wrong and that they are fixing their self made wrong mistake. This analogy is therefore not accurate as error rate can actually decrease over time.
- kykat 1y agoIf we assume that each action has 99% success rate, and when it fails, it has 20% chance of recovery, and if the math here by gemini 2.5 pro is correct, that means the system will tend towards 95% chance of success. === In equilibrium, the probability of leaving the Success state must equal the probability of entering it. (Probability of being in S) * (Chance of leaving S) = (Probability of being in F) * (Chance of leaving F) Let P(S) be the probability of being in Success and P(F) be the probability of being in Failure. P(S) * 0.01 = P(F) * 0.20 Since P(S) + P(F) = 1, we can say P(F) = 1 - P(S). Substituting that in: P(S) * 0.01 = (1 - P(S)) * 0.20 0.01 * P(S) = 0.20 - 0.20 * P(S) 0.21 * P(S) = 0.20 P(S) = 0.20 / 0.21 ≈ 0.95238
- ninetyninenine 1y agoThat’s math based off of arbitrary initial assumptions. There are numbers that work. All this math is useless. Use your brain. The entire point I’m communicating is that it’s not a given that it must become less accurate. There are multiple open possibilities here and scenarios that can occur. Doing random math here as if you’re dropping the mic is just pointless. It doesn’t do anything. It’s like making up a cosmological constant and saying the universe is collapsing look at my math.
- darkwater 1y agoI saw them too. And after that, slip in another mistake.